darkknight:

Loudspeaker manufacturer advertises that their model no. 801 speaker produces a sound pressure level of 87 decibels (db) when a reference test tone is applied. A competing speaker company advertises that their model X-1 speaker produces a sound pressure level of 25 db when fed the same test signal. What is the ratio of the two sound intensities produced by these speakers? If you wanted to find a speaker which produces a sound intensity thrice that of no.801 when fed the test signal, what is its sound pressure

darkknight:

The formula for sound pressure of a noise is defined by B = (10)(log (I/Io)) where I is the intensity of sound in watts per meters squared and Io is hearing threshold for a human, lowest sound human can hear: 10^(-12) watts per meters squared

darkknight:

what do I do?

DuarteME:

Let me just rewrite everything here so that you can see everything at once: - We have \(B_{801} = 87\textrm{ dB}\) and \(B_{\textrm{X}-1} = 25\textrm{ dB}\); - We know that \(B = 10 \log_{10}\left(\dfrac{I}{I_0}\right)\), with \(I_0 = 10^{-12}\textrm{ W m}^{-2}\); - We want to compute \(\dfrac{I_{801}}{I_{\textrm{X}-1}}\). You can start by expressing \(I\) in terms of \(B\) to compute \(I_{801}\) and \(I_{\textrm{X}-1}\) first and then take their ratio. If you notice that \(B\) is defined in terms of \(\log_{10}\), you can also take the difference \(B_{801} - B_{\textrm{X}-1}\) and use the properties of the logarithm to extract the ratio directly. Can you see how? (:

Imsmart:

Nice u stupe

darkknight:

is the answer to What is the ratio of the two sound intensities produced by these speakers? 10^6.2?

DuarteME:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight is the answer to What is the ratio of the two sound intensities produced by these speakers? 10^6.2? \(\color{#0cbb34}{\text{End of Quote}}\) That is correct: \(\dfrac{I_{801}}{I_{\textrm{X}-1}} = 10^{\frac{B_{801} - B_{\textrm{X}-1}}{10}} = 10^{6.2} \quad \ddot{\smile}\) Can you now answer the next one?

darkknight:

oh right

darkknight:

give me a couple min

darkknight:

is 3 times the sound of intensity of no. 801 (which is 10^-3.3) 10^-9.9 or 10^-1.1?

darkknight:

or 3(10^-3.3)

DuarteME:

The sound intensity of 801 is given by: \(I_{801} = I_0 10^{B_{801}/10} = 10^{-12} \times 10^{87/10} \textrm{ W m}^{-2} = 10^{-3.3}\textrm{ W m}^{-2}.\) Therefore, thrice this sound intensity is: \(I = 3 \times I_{801} = 3 \times 10^{-3.3}\textrm{ W m}^{-2}.\)

darkknight:

k thats what I thought

darkknight:

about 9.177

darkknight:

right?

DuarteME:

What is 9.177 supposed to be? Notice that the 'sound pressure' of the 801 speaker is 87 dB and the new speaker is supposed to have a larger intensity.